OpenStudy (anonymous):

Compute the modulus and argument of each complex number. a) -5i b) sqrt 10(cos(5pi/7) + i sin (5pi/7))

OpenStudy (anonymous):

I got the modulus. What's the argument?

OpenStudy (anonymous):

a) modulus 5,arg \[-0.5\pi\] b) modulus \[\sqrt{10}\],arg \[5\pi/7\]

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

u r wlcm :)